A man distributes £2 among 20 boys,giving some 5p and 25p to the rest. How many boys receive 25p each.

Answer by ikleyn(52787)

.
25*x + 5*(20-x) = 200


25x + 100 - 5x = 200


25x - 5x = 200 - 100


20x = 100    ======>  x = 100/20 = 5.


ANSWER.  5 lucky boys get 25 p each.